OpenStudy (anonymous):

Verify the identity. cot (x-pi/2)= -tan x

OpenStudy (anonymous):

Working on the left hand side, you get: \[\frac{ 1 }{ \tan (x-\frac{ \Pi }{ 2 }) }\] After that, follow the property that:\[\tan(A-B)=\frac{ tanA-tanB }{ 1+tanAtanB }\] Giving us: \[\frac{ 1 }{ \frac{ tanx-\tan \frac{ \Pi }{ 2 } }{ 1+tanxtan \frac{ \Pi }{ 2 } } }=\frac{ 1+tanxtan \frac{ \Pi }{ 2 } }{ tanx-\tan \frac{ \Pi }{ 2 } }\]

OpenStudy (anonymous):

From here, we can use the calculator to find tan pi/2 .

hartnn (hartnn):

or you could convert cot into cos/sin and directly use, cos (x-pi/2) = sin x sin (x-pi/2) = -cos x